Eoma Einancial Statements Learning.man Canital Budgeting eting Question 1 If a project is expected to cost $150,000 and produce cash flows over the next four years of $75,000, $0, $55,000 and $60,000 respectively, what is the discounted payback, assuming a discount rate of 7.25%? 3.33 3.78 4.00 2.33 Submit answer O Exit
Answer #1

Cashflow - CF units in $

CF yr 0 = -150000

CF yr 1 = 75000; discounted present value = 75000/(1+7.25%)=69930.07

CF yr 2= 0; discounted present value = 0

CF yr 3= 55000; discounted present value = 55000/(1+7.25%)^3=44583.15

CF yr 4 = 75000; discounted present value = 60000/(1+7.25%)^4=45348.41

Sum of cashflow till year 3 =-150000+69930.07+0+44583.15=-35486.78

In year 4, we need 35486.78/45348.41 of a year to get to a cumulative value of 0 which is the payback = 0.78 of 4th year; so a total of 3.78 years reqd.